Ferric chloride, FeCl, reacts with silver nitrate, AgNO3 to form ferric nitrate, Fe(NO3)3 andsilver chloride, AgCl.a. Write the balanced chemical equation of the reaction.b. If 108.5 moles of FaCl, is combined with 76.8 moles of AgNO3, how many moles ofAgCl is produced?c. How many grams of AgCl is produced?d. Which reactant is the limiting reagent?e. Which reactant is the excess reagent?​

1
Expert's answer
2021-05-18T17:52:56-0400

"(a)3AgNO_3+FeCl_3\to Fe(NO_3)_3+3AgCl"


  

(b) 108.5 mole of "FeCl_3" react with 76.5 mole of "AgNO_3"


    "25.6 mole FeCl_3 + (25.6\times 3AgNO_3)_4" "\to" (25.6"\times" 3) AgCl



(b) 108.5 mole of "FeCl_3" react with 76.5 mole of "AgNO_3"


   "25.6 \text{ mole } FeCl_3 + (25.6\times 3) AgN)_4\to (25.6\times 3) AgCl"


   So, 76.8 mole AgCl produced and (108.5-25.6) mole=82.9 mole "FeCl_3"


(c) 1mole AgCl=143.32 mol


  76.8 mole AgCl = 1006.98 gm AgCl produced.


(d) AgN)"_3" completely reacted as finished in the mixture, So "AgNO_3" is limiting reagent.


(e) "FeCl_3" is excess reagent.